OpenStudy (moongazer):

What is the root of -sqrt of x-5 = 1 - sqrt of x

OpenStudy (anonymous):

first, square both sides then you will get x-5=1-2sqrt of x+x x-x-5-1=-2sqrt of x -6=-2sqrt of x sqrt of x = 3, then square both sides, you will get x=9.